### v1

# SelectLLM: Query-Aware Efficient Selection Algorithm for Large Language Models

Abstract
Large language models (LLMs) have gained increased popularity due to their remarkable success across various tasks, which has led to the active development of a large set of diverse LLMs. However, individual LLMs have limitations when applied to complex tasks because of such factors as training biases, model sizes, and the datasets used. A promising approach is to efficiently harness the diverse capabilities of LLMs to overcome these individual limitations. Towards this goal, we introduce a novel LLM selection algorithm called SelectLLM. This algorithm directs input queries to the most suitable subset of LLMs from a large pool, ensuring they collectively provide the correct response efficiently. SelectLLM uses a multi-label classifier, utilizing the classifier’s predictions and confidence scores to design optimal policies for selecting an optimal, query-aware, and lightweight subset of LLMs. Our findings show that the proposed model outperforms individual LLMs and achieves competitive performance compared to similarly sized, computationally expensive top-performing LLM subsets. Specifically, with a similarly sized top-performing LLM subset, we achieve a significant reduction in latency on two standard reasoning benchmarks: 13% lower latency for GSM8K and 70% lower latency for MMLU. Additionally, we conduct comprehensive analyses and ablation studies, which validate the robustness of the proposed model.
SelectLLM: Query-Aware Efficient Selection Algorithm
for Large Language Models
Kaushal Kumar Mauryathanks: Equal contribution                        KV Aditya Srivatsa11footnotemark: 1 Mohamed Bin Zayed University of Artificial Intelligence, Abu Dhabi, UAE {kaushal.maurya, vaibhav.kuchibhotla, ekaterina.kochmar}@mbzuai.ac.ae                        Ekaterina Kochmar
Figure 1: Accuracy with majority vote accuracy for the test sets of GSM8K and MMLU against the number of LLMs considered from a diverse model pool. Models are added as per descending order of their performance on corresponding training sets.
1 Introduction
In recent years, large language models (LLMs) have demonstrated remarkable capabilities in solving a wide range of core NLP tasks (Bommasani et al., 2021; Chang et al., 2023). Despite these advancements, it has been noted that existing LLMs struggle with complex tasks such as factually-grounded reasoning and planning (Wei et al., 2022; Kojima et al., 2022; Minaee et al., 2024). Moreover, the wide range of LLMs available seem to exhibit diverse capabilities (Jiang et al., 2023), resulting in no single (especially open-source) LLM being effective across all benchmarks and datasets.
Though newer and larger models are being trained from scratch to close this gap, an alternative and cost-effective approach involves harnessing the diverse capabilities of existing models to improve the overall response quality using ensembling (Wang et al., 2022, 2023; Li et al., 2024) and collaborative frameworks (Wu et al., 2023b; Li et al., 2023). However, these approaches often require access to the responses from all models in the pool to choose the optimal response(s), which greatly increases the overall computational cost for such ensembles.
Keeping in mind the diverse capabilities of LLMs, not all models may be apt for all kinds of tasks. Figure 1 reports the majority-vote accuracy of a model pool (spanning up to 7 diverse LLMs) on two challenging reasoning benchmarks – GSM8K (Cobbe et al., 2021) and MMLU (Hendrycks et al., 2021). As the graph demonstrates, utilizing more LLMs initially improves performance, which is supported by previous research towards employing more LLMs and more responses per model (Li et al., 2024). However, note that using more (or even all) models in the pool does not necessarily result in the best scores. Thus, selectively abstaining from querying unsuitable LLMs for a given task may help improve the overall response quality of such ensembles. Additionally, such an approach would implicitly save computation by accessing fewer models per query.
In this paper, we propose the SelectLLM algorithm to explore this idea. Our approach first learns the model-specific capabilities for different LLMs through a dataset of diverse queries using a classifier. When performing inference for an unseen query, this insight is used to predict confidence scores for each LLM in the model pool, representing how likely they are to solve the underlying task. We create and test multiple selection policies that determine the optimal set of LLMs for the given query based on their respective confidence scores.
Our findings indicate that the proposed system: (i) Outperforms individual LLMs in terms of their accuracy and (ii) Achieves competitive accuracy score with a similar sized top-performing subset while significantly reducing latency costs. The contributions of our work are as follows:
  • We introduce the novel SelectLLM algorithm, which is based on a multi-label classifier and an optimal confidence-based policy. This approach efficiently navigates input queries to the ideal subset of LLMs from a larger pool to improve response quality and also reduce computational costs in the process. To the best of our knowledge, this is the first study to propose LLM selection method.
  • The efficacy of the proposed algorithm is evaluated on two challenging reasoning benchmarks. We report a gain of $6.67$ points on GSM8K and $2.08$ points on MMLU in terms of accuracy compared to the highest scoring individual LLMs in the corresponding model pools. We also report significantly lower latencies (13% for GSM8K and 70% for MMLU) compared to similar sized top-performing LLM subset from the pools.
  • Through empirical analysis and ablation studies across two datasets and 7 LLMs, we demonstrate that the proposed model is reliable, robust, and cost-efficient.
Figure 2: Overview of the proposed workflow.
LLM GSM8K MMLU
llama2-7b-lm 23.24 47.87
gemma-7b-lm 73.13 66.44
mistral-7b-lm 59.41 62.34
metamath-7b-lm 88.57 42.88
*gemma-7b-it 41.73 50.57
llama2-13b-chat 50.19 54.25
*mistral-7b-it 55.96 53.48
Table 1: Accuracy with majority voting (maj@10) for considered LLMs on GSM8K and MMLU datasets with all three splits. All scores were calculated over 10 response generations for each LLM. * The term ‘it’ indicates instruction-tuned LLMs.
2 Related Work
Model Diversity
Recent surveys (Bommasani et al., 2021; Minaee et al., 2024) suggest that LLMs can develop emergent capabilities. Specifically, this means that models can show behavior and demonstrate skills beyond those explicitly taught. By virtue of changing the training data, models can also be trained to exhibit a wide variety of domain expertise. At the same time, Jiang et al. (2023) demonstrate that no single open-source LLM outperforms other models across popular benchmarks. This further motivates the need to develop ensembling methods to improve the combined performance of a pool of LLMs with diverse capabilities.
Model Selection
Routing queries among LLMs involves aligning their capabilities with the underlying tasks from the input queries. However, selecting models for routing with LLMs differs from that in traditional ML (Bishop, 2006; Raschka, 2020) due to the disparity between the training and test datasets. LLMs are trained on massive corpora with straightforward objectives like next-token prediction (Brown et al., 2020), while test data often includes complex tasks like reasoning and question answering (Hendrycks et al., 2021; Cobbe et al., 2021; Joshi et al., 2017), summarization (Tam et al., 2023), and classification (Zhang et al., 2023), which may not be prevalent in the training data. This discrepancy makes it challenging to assess the difficulties in resolving complex queries. Additionally, studies like  Rabinovich et al. (2023) and  Srivatsa and Kochmar (2024) suggest that prompt characteristics such as length and readability can significantly impact an LLM’s ability to handle tasks.
LLM Ensembling
Previous attempts at ensembling and routing LLMs typically fall into two categories: (1) Selecting the best response from multiple LLM generations, as seen in Liu and Liu (2021), Ravaut et al. (2022), and Jiang et al. (2023). However, this approach requires querying all LLMs in the model pool for each query during inference, which can be computationally expensive with a large number of LLMs. (2) Minimizing the number of queries to larger LLMs to reduce latency and computational costs, as demonstrated by Shnitzer et al. (2023) and Ding et al. (2024), who redirect simpler queries to the smallest model capable of handling the task to minimize querying costs with minimal drop in performance. Tackling both tasks by routing to the single-best LLM has shown to be challenging (Srivatsa et al., 2024). In this paper, however, we aim to develop an algorithm that can improve response accuracy above individual LLMs and their combinations by querying only the subset of LLMs expected to be capable of solving the input query. This in turn, saves computational costs and reduces latency by avoiding unnecessary queries made to unsuitable LLMs.
3 Problem Setting
We propose an ensembling-based LLM inference algorithm – SelectLLM – to efficiently select the most suitable query-aware few LLMs from a large pool of available LLMs. The algorithm harnesses the diverse capabilities of different LLMs and selects a subset of models for the input query, jointly leading to the correct response and this selection of small subset leads to the reduction in latency.
Formally, for a given set of input queries $Q=\{q_{1},q_{2},\dots,q_{n}\}$ and a pool of LLMs $L=\{l_{1},l_{2},\dots,l_{k}\}$, the objective is to learn a selection model $M_{r}\in\{0,1\}^{k}$ (where 1 means that $L_{k}$ can answer the query correctly) such that each user query $q_{i}\in Q$ is directed to a subset $L_{s}$ with latency lower than the whole pool (i.e., $latency(L_{s})<latency(L)$). Here, $m_{i}$ is the $i^{th}$ predicted label by $M_{r}$, $L_{i}$ is the corresponding $i^{th}$ LLM in $L$, and $latency(L_{s})$ is the sum of the individual LLM latencies in $L_{s}$. $L_{s}$ consists of the top-$s$ LLMs based on the confidence scores predicted by $M_{r}$: in the best-case scenario, a query $q_{i}$ is processed by a single LLM, while in the worst-case scenario, $q_{i}$ is processed by all LLMs in $L$. We find that this algorithm favors optimal performance, surpassing individual LLM capabilities by a significant margin. Furthermore, it achieves competitive accuracy compared to top-performing similar size LLM subsets and maintaining lower latency.
4 Methodology
This section discusses the methodology behind LLM sampling and formally introduces the SelectLLM algorithm.
4.1 LLM Sampling
4.1.1 Selection of Benchmarks and LLMs
Recent studies suggest that LLMs struggle with reasoning and mathematical tasks (Patel et al., 2021; Wu et al., 2023a). This study focuses on two challenging reasoning tasks: specifically, we use the GSM8K dataset by Cobbe et al. (2021) for mathematical reasoning and the MMLU dataset by Hendrycks et al. (2021) for natural language understanding/reasoning tasks. GSM8K consists of 8,792 diverse grade-school level math word problems (MWPs) in English, while MMLU contains 15k multiple-choice questions in English, spanning 57 subjects across STEM, humanities, social sciences, etc. The dataset statistics for all three splits are summarized in Appendix Table 5.
We have selected a diverse set of LLMs based on explicit and implicit criteria. The explicit criteria encompass performance on benchmarks, training methodologies, model specialization, and modes of operation (chat and non-chat), among others. Some of these diverse attributes are presented in the Appendix Table 5. The implicit criteria include factors such as diverse inference latencies (refer to Table 6) and prompting types (i.e., zero-shot vs. few-shot), among others. Further, we consider relatively small open-source LLMs because: (i) Experiments with these LLMs are suitable for an academic lab setup; (ii) This aligns with the research trend towards developing LLMs suitable for small mobile devices Abdin et al. (2024); and (iii) We hypothesize that the proposed modeling approach, which is LLM-agnostic, will work for smaller LLMs and be scalable to larger LLMs, although we leave the proof of this to future work.
Based on the recent findings about the appropriate usage of prompts (Sahoo et al., 2024) and from our own experiments, we have tailored few-shot chain-of-thought (COT) prompts for non-chat LLMs and zero-shot COT prompts for chat LLMs. More details on LLM sampling and answer extraction are available in the Appendix Section A.
4.1.2 Data Preparation for the SelectLLM Model: SLData
In this study, we evaluate the performance of each LLM by generating 10 responses for each input query to ensure reliable and replicable behavior of the proposed model. We employ Majority Voting Li et al. (2024) to assess whether a query is correctly answered by the LLM or not. Specifically, Majority Voting (maj@k $\in\{0,1\}$) determines whether the most frequent answer from an LLM matches the gold answer or not. The accuracy with maj@10 across all input prompts is reported in Table 1. In the rest of this paper, we consider only those LLMs for which the answer extracted viability scores are above 90% (see more details in Appendix Section A) to ensure response reliability, resulting in 6 acceptable LLMs for the GSM8K dataset and 7 for the MMLU dataset. We prepare the training dataset for SelectLLM by associating each input query with the LLM(s) whose majority vote answer (across 10 samples) matches the gold answer, i.e., maj@10 $=1$. Formally, the target label for a query prompt $q\in Q$ is given by $label\left(q\right)=\left\{l\>|\>l\in L,maj@10\left(q,l\right)=1\right\}$, where $L$ is the set of candidate LLMs and $Q$ is the set of question prompts from GSM8K or MMLU. We denote this dataset as SLData.
Algorithm 1 SelectLLM Inference Algorithm with WeightedMaxConf Policy
1:Queries $\mathcal{Q}$, LLMs $\mathcal{L}$, SLData $(\mathcal{X},\mathcal{Y})\in\mathcal{D}$, pre-trained model $\mathcal{M}_{\theta}$ with parameters $\theta$
[…]
3:for each query $q_{i}$ in $\mathcal{Q}$ do
4:     Perform inference: $q_{i}^{logits},q_{i}^{labels}=\mathcal{M}_{\hat{\theta}}(q_{i})$triangleright Using fine-tuned model $\mathcal{M}_{\hat{\theta}}$
5:     Calculate confidence scores: $c_{i}=\sigma(q_{i}^{logits})$triangleright $\sigma$ is the sigmoid activation function
6:     Select top-$s$ confidence scores: $c_{i}^{(s)}=\max(c_{i},s)$ and associated LLMs: $L_{s}=L(c_{i}^{(s)})$
7:     Initialize an empty set for answer set $A_{i}\leftarrow\phi$
8:     for each $l_{j}$ in $L_{s}$ do
9:         Generate 10 responses: $a_{i}^{10}=l_{j}(q_{i})$triangleright Generate 10 responses with LLM $l_{j}$
10:         Find answer frequency $a_{i}^{f}=\{a_{i}^{k}:\text{countof}(a_{i}^{k})/\sqrt{c_{i}^{(s)j}}\,|\,\text{for }a_{i}^{k}\text{ in unique}(a_{i}^{10})\}$
11:         $A_{i}\leftarrow A_{i}\cup a_{i}^{f}$
12:     end for
13:     Return most frequent answer in $A_{i}$
14:end for
4.2 The Proposed SelectLLM Algorithm
Prior research (Ding et al., 2024) indicates that easy queries are correctly solved by smaller or general-purpose LLMs, whereas complex queries necessitate the use of specialized or larger LLMs. Conversely, there are rare queries that are incorrectly responded to by large LLMs but correctly responded to by small LLMs Nezhurina et al. (2024). Due to a lack of widely established query-to-LLM mappings, brute-force approaches are typically employed, querying every available LLM to obtain correct answers. As language models continue to advance rapidly, resulting in a large number of LLMs, such approaches become computationally inefficient and sometimes infeasible.
To address this challenge, it is essential to explore modeling approaches to harness the capabilities of different LLMs (i.e., the ability to respond correctly to input queries) jointly and efficiently. One promising approach involves identifying and directing input queries to the most suitable subset of LLMs from the large pool. This ensures that the query is accurately addressed while maintaining lower latency compared to running inference on all LLMs in the pool. Towards this objective, we introduce the SelectLLM algorithm, designed to select a tailored subset of LLMs, taking into account the nature of the query and enabling them to collaboratively provide correct responses efficiently. The SelectLLM algorithm comprises two primary components: (i) A multi-label classifier (MLC) which is fine-tuned with SLData dataset, and (ii) A selection policy, which utilizes MLC’s prediction confidence scores (i.e., the likelihood of an LLM responding correctly to the query) to determine the suitable subset of LLMs.
Multi-label Classifier (MLC)
We fine-tune the pre-trained RoBERTa model (Liu et al., 2019) using SLData with a multi-label classification objective. The fine-tuned model predicts a set of LLMs (i.e., LLMs’ identities) that are capable of solving the input query. Specifically, the model learns to predict the most suitable LLMs along with corresponding confidence scores $C$. Since the multi-label classification (MLC) model is based on an encoder-only architecture, the cost of MLC inference is negligible compared to running inference with large auto-regressive language models (Sun et al., 2019). We have achieved weighted F1 scores of 0.71 and 0.68 for the GSM8K and MMLU test datasets, respectively for MLC.
Confidence-Based Policies
In this section, we discuss how the confidence scores $C$ are utilized to select a suitable subset of LLMs $L_{s}$ for each input query. $C$ is defined as $\{c_{1},c_{2},\dots,c_{i},\dots c_{n}\}$ where $c_{i}$ is the confidence scores for the $i^{th}$ input query. Each $c_{i}$ is further represented as $\{c_{i}^{l_{1}},c_{i}^{l_{2}},\dots c_{i}^{l_{j}}\dots c_{i}^{l_{k}}\}$ where $c_{i}^{l_{j}}$ is the confidence score of the $j^{th}$ LLM for the $i^{th}$ query.
The performance of SelectLLM is determined by the selection policy used. For example, a greedy policy that always selects the LLM (or set of LLMs) with the highest confidence can be suboptimal. This is because another LLM (or set of LLMs) in the pool might have higher accuracy but may not be chosen due to slightly lower confidence. Additionally, when there are two subsets with similar confidence and accuracy, it is more efficient to select the subset with lower cumulative latency. Considering these aspects, we propose the following three optimal policies:
  1. 1.
    LabelledMaxConf: This policy selects the top-$s$ LLMs ($L_{s}$) for an input query $q_{i}$ based on two constraints: (i) the LLMs should be present in the MLC predictions, and (ii) Only those LLMs that have confidence scores within the top-$s$ from $c_{i}$ are considered.
  2. 2.
    MaxConf: This is a more flexible policy than LabelledMaxConf as it only takes into account the second constraint, i.e., it selects the top-$s$ LLMs corresponding to the top-$s$ confidence scores from $c_{i}$.
  3. 3.
    WeightedMaxConf: This policy begins by selecting the top-$s$ LLMs based on their high confidence scores, i.e., for a given query $q_{i}$, we denote the selected LLMs as $L_{s}^{q_{i}}$. Subsequently, we modify the frequency of answer values extracted from the responses of each selected LLM, which involves dividing the frequency of each value by the square root of the confidence score associated with the respective LLM. Finally, we collect all response values and their modified frequencies across the selected LLMs (frequencies are added for the same value). The value with the highest frequency after the modification is selected as the final response. The formal steps are presented in Algorithm 1. Intuitively, dividing by the square root of the confidence score aims to mitigate biased predictions by the classifier. This adjustment ensures fairer opportunities for each selected LLM to contribute to the majority voting process.
Across all three policies, in case of a conflict where two LLMs have similar confidence, the light-weight LLM (i.e., with lower latency) is preferred.
5 Experimental Setup
5.1 Baseline Models
The following baseline models are included for comparison:
  1. 1.
    Oracle: The maximum performance is assumed under the premise that an oracle always selects the lowest latency subset of LLMs that generates the correct majority vote answer for each question (if possible; otherwise, the question attempt is marked as incorrect). Empirically, this is obtained by evaluating all subsets of LLMs, i.e., ($12^{k}-$ 1), where $k$ is the total number of LLMs.
  2. 2.
    Random: This represents the mean performance of uniformly randomly selecting an LLM subset from all possible ($12^{k}-$ 1) subsets for each query. We report mean evaluation scores across 1,000 independent runs to avoid biases.
  3. 3.
    Individual Models: This is the performance of individual models with maj@10.
  4. 4.
    All LLMs: This baseline reports the mean accuracy of maj@(10×$\times|L|$) based on the combined pool of 10 generations from each LLM and is similar to the approach of Li et al. (2024).
  5. 5.
    LLM-Blender Jiang et al. (2023): An ensembling framework was developed to utilize the diverse strengths of multiple open-source LLMs. Specifically, it employs PairRanker, which utilizes a cross-attention-based method for pairwise comparison of different LLM responses to determine the superior one. We use the officially released model checkpoint in our setting.
  6. 6.
    Top-s LLMs: For this baseline, we consider the responses of the $s$ top-scoring LLMs for the majority-vote. We determine the top performing models by the aggregate accuracies of each model on the test sets of corresponding datasets.
For All LLMs, LLM-Blender and Top-s LLMs, the latency remains constant since it necessitates inference with all LLMs to determine the performance.
5.2 Evaluation Metrics
We evaluate the performance of all models with the accuracy (Acc) metric using majority voting (see Section 4.1.2). Additionally, we report the latency per query (Lat) to estimate efficiency. The exact costs of model execution, including factors like latency, FLOPs, and energy consumption, may vary and are influenced by factors such as prompt templates, hardware capabilities, and network connectivity, especially in LLM inference scenarios. To ensure a fair comparison, we record the inference latency of each LLM under uniform conditions using single A100 GPUs. The individual latencies for each LLM are detailed in Appendix Table 6.
6 Results and Discussion
Models / Setups GSM8K MMLU
Acc ($\uparrow$) Lat ($\downarrow$) Acc ($\uparrow$) Lat ($\downarrow$)
Oracle 90.52 3.24 90.46 1.75
Baseline Random 69.49 9.65 58.20 8.27
All LLMs (Li et al., 2024) 76.04 19.00 60.92 16.40
LLM-Blender Jiang et al. (2023) 75.28 19.00 60.27 16.40
Top-s LLMs 77.48 (s=4) 19.00 65.75 (s=2) 16.40
gemma-7b-lm 71.27 7.10 63.73 3.00
mistral-7b-lm 60.50 3.70 61.57 1.80
metamath-7b-lm 67.25 4.70 41.76 2.40
llama2-7b-lm 48.10 2.30
llama2-13b-chat 49.20 1.80 52.94 4.80
mistral-7b-it 56.71 1.00 53.92 1.10
gemma-7b-it 42.23 0.70 50.72 1.00
SelectLLM MLC + LabelledMaxConf 75.66 (s=4) 14.69 65.68 (s=2) 4.78
MLC + MaxConf 77.48 (s=4) 16.50 65.68 (s=2) 4.78
MLC + WeightedMaxConf 77.94 (s=4) 16.50 65.81 (s=2) 4.78
[…]
Table 2 presents the performance results for the oracle, baselines, and the proposed SelectLLM models across both GSM8K and MMLU datasets. We have also reported respective inference latencies to analyze the efficiency of different models. We make the following major observations.
Performance of Individual Models: Among the considered LLMs, metamath-7b-lm, a specialized mathematical LLM, emerges as the best performing model for the mathematical reasoning task (GSM8K) but performs worse on the natural language understanding/reasoning task (MMLU). This indicates the LLM diversity and the need for the effective combination of their capabilities. Gemma-7b-lm performs the best on MMLU and second best on the GSM8K data. Chat LLMs perform poorly on GSM8K; however, for MMLU, their performance is mixed; some non-chat LLMs achieve better scores. The latency for non-chat models is expected to be higher than for chat models, as they utilize a 5-shot COT, whereas chat models employ a 0-shot COT strategy.
Performance of Baselines: We have tested four baseline models (Random, All LLMs, LLM Bender, and Top-s LLMs) to understand the effect of ensembling LLMs. It can be observed that the performance of the Random baseline model is better than that of many individual LLMs, demonstrating the promise of utilizing multiple LLMs. To our surprise, including all LLMs does not perform better than selecting only the top few. This contradicts the findings of Li et al. (2024) and can be attributed to two reasons: (i) the authors have used only a few LLMs, and (ii) they have employed very powerful LLMs, leaving out weaker LLMs. The Top-s LLMs model emerges as the best baseline, showing the promise of selecting the best-performing LLMs. However, the latency of all these baselines is high (except Random baseline where accuracy is lower), making them impractical for real-world usage.
The Effect of Different Policies with the SelectLLM Algorithm: It can be observed that with the policy LabelledMaxConf, the performance on the GSM8K dataset is the lowest. This may be because the policy relies on both the MLC prediction and the confidence scores, where the classifier almost always predicts metamath-7b-lm as the class label (because metamath-7b-lm is a specialized model for math, and 88% of the SLData training has this label). However, this impact is negligible for MMLU, where the label distribution for all LLMs is not as highly skewed. This limitation is overcome by MaxConf and WeightedMaxConf, which relax the constraint on MLC label prediction and only operate on confidence scores. This allows models to incorporate other LLMs and push the performance, particularly for the GSM8K data. Mathematical tuning in WeightedMaxConf allows policies to select LLMs more effectively and improve the scores. Overall, the WeightedMaxConf policy emerged as the best performing, with a slight edge over its close competitor, the MaxConf policy.
SelectLLM vs. Baseline Models: The proposed SelectLLM with WeightedMaxConf outperforms individual LLMs baseline by a large margin across both datasets. This indicates that inference with more LLMs boosts performance with a cost of slightly higher latency. Furthermore, the model performs better than the Random, All LLMs and LLM-Blender while being competitive with the Top-$s$ LLMs baseline with respect to accuracy. At the same time, the proposed model has much lower latency (13% lower for GSM8K and 70% lower for MMLU) than Top-$s$ LLMs, indicating the efficiency of the proposed model. This shows that SelectLLM achieves the set goal of better performance than individual LLMs combined with efficiency.
Performance of Oracle vs. other Models: It can be observed that the Oracle score is much higher than that of any individual LLM, baselines, and SelectLLM. This indicates a considerable scope for improvement in the LLM ensembling.
Figure 3: Latency vs. accuracy vs. $s$-values plots for different SelectLLM policies on GSM8K (left) and MMLU (right) datasets. $s$-value: total number of LLMs selected with SelectLLM.
Latency vs. Accuracy vs. $s$-Value: Figure 3 shows the relationship between latencies, accuracies, and $s$-values for different SelectLLM policies across both datasets. WeightedMaxConf consistently outperforms MaxConf and WeightedMaxConf on both accuracy and latency for most $s$-values (except 5 and 6 for GSM8K). This indicates the superiority of WeightedMaxConf, maintaining effectiveness even with a small number of LLMs. A larger number of LLMs and higher latency are required for GSM8K, while a lower $s$-value and latency are needed for MMLU to achieve high accuracy. Selecting 4 LLMs for GSM8K and 2 for MMLU achieves optimal performance, indicating the effectiveness and efficiency of SelectLLM with WeightedMaxConf.
Query Awareness Analysis of SelectLLM: With this analysis, we aim to understand how the distribution of selected LLMs changes as more LLMs are selected (with increasing values of $s$) in SelectLLM algorithm. Figure 5 in the Appendix presents such distribution. For both datasets, in the top-1 and top-2 subsets, most of the queries are directed to the best-performing LLMs. However, as the subset size increases, the dominance of the top-performing models diminishes, leading to a more uniform distribution where queries are routed towards more LLMs to boost the performance. This indicates the input query awareness of the SelectLLM model, which is adept at assigning a suitable set of LLMs for the input query.
Prediction Distribution Analysis: Appendix Figure 6 presents the distribution of the number of input queries with correct answer predictions using the top-3 individual LLMs and models with the SelectLLM algorithm across both datasets. It can be observed (count from the right column) that the proposed model is able to correctly provide answers to input queries compared to other individual LLMs, which supports the performance gain reported in Table 2. The distribution also indicates (count from the bottom) that the proposed model utilizes the capabilities of multiple individual LLMs to extract the correct answers. Additionally, Figure 8 illustrates the subject-wise performance differences between SelectLLM and the best performing individual LLM (i.e., gemma-7b-lm) for MMLU. It can be observed that the proposed model shows substantial gains in the majority of subjects, while performing slightly worse for a few of them, showing scope for improvement.
Upper-bound Performance of the SelectLLM: In this section, we investigate potential reasons for the performance gap between the proposed model and the Oracle. With this aim, we estimate the upper-bound performance of SelectLLM with WeightedMaxConf policy. Specifically, we assess all subsets of the top-$s$ LLMs, predicted with SelectLLM for each $q_{i}$. A query is considered solved if at least one subset yields a correct answer. Table 3 presents Oracle, SelectLLM upper bound, and SelectLLM scores across both datasets. The results reveal the following insights: (i) the classifier’s performance is constrained (weighted F1 score for GSM8K is 0.71 and for MMLU is 0.68) due to limited training data (approximately 7K for GSM8K and 14K for MMLU), which likely contributes to the performance gap. Therefore, augmenting the training data or improving the classifier model could enhance scores. (ii) A performance disparity between the best policy and the upper bound suggests the potential for developing better policies (specifically for MMLU). However, since policies rely on the classifier’s confidence scores, enhancing the classifier could also bridge this gap.
Models GSM8K MMLU
Oracle 90.52 90.46
Upper Bound 78.77 76.20
SelectLLM 77.94 65.81
Table 3: Accuracy in different experimental setup.
Ablation Studies: Does the proposed algorithm is effective with different LLMs pool sizes (i.e., value of $k$)? To investigate this question, we conduct ablation studies considering various LLM pool sizes, i.e., $k={1,...,6}$, for the GSM8K dataset. We examine two extreme settings: pool with top-$k$ and bottom-$k$ LLMs based on individual LLM performance. This encompasses many configurations as LLMs with similar or different performances may be present in the LLM pool. We also compare this with closet strong Top-$s$ LLM baseline. The results are shown in the Appendix Figure 7. We observe that even with different $k$ values across both top-$k$ and bottom-$k$ setups, the proposed SelectLLM outperforms (in terms of accuracy) the Top-$s$ LLM baseline. This indicates that the proposed approach is LLM pool-agnostic. Moreover, as the number of $k$ values increases, latency becomes a factor: for larger pool sizes, the latency for SelectLLM is much lower than for Top-$s$ LLM. Similar results are observed with the MMLU dataset.
7 Conclusions and Future Directions
To the best of our knowledge, this paper presents the first study to efficiently navigate input queries to the most suitable subset of LLMs selected from a large pool. We introduce the novel SelectLLM algorithm, which selects a lightweight subset of LLMs using a multi-label classifier and confidence-based optimal policies. The model is evaluated with two challenging reasoning datasets and compared with several strong baseline models. SelectLLM outperforms individual LLMs and achieves competitive accuracy with a similar subset size of the top-performing LLMs while maintaining significantly lower latency (13% lower latency for GSM8K and 70% lower for MMLU). Although the proposed modeling is promising, we believe better modeling can push the performance closer to the oracle performance. Our findings serve as a strong foundation for such future modeling endeavors. A feasible direction could involve incorporating LLM-related and question-related features to enhance query and LLM awareness in modeling.
Limitations
One of the key limitations of the proposed SelectLLM algorithm is the limited availability of training data for the multi-label classifier, with only 7K instances for GSM8K and 14K for MMLU. This limitation can potentially lead to biased learning. Despite several measures to address this issue, such as weighing labels to counteract label imbalance, conducting extensive optimal hyperparameter searches, experimenting with different sizes of probabilistic and LLM-based models (with RoBERTa performing the best), and obtaining the best checkpoint with the validation set, the performance remains suboptimal. The algorithm achieves a weighted F1 score of 0.71 for GSM8K and 0.68 for MMLU. Additionally, we could only extract viable answers for 83% to 95% of queries using different LLMs. For the remaining queries, the answers extracted by our algorithm may be invalid or incorrect. These invalid answers could be due to limitations in the extraction algorithm or the LLM itself, where the LLM fails to provide answers in an extractable format.
Ethics Statement
This paper introduces a novel SelectLLM algorithm to effectively harness the power of LLMs with different capabilities. As the proposed routing models use LLMs, we must acknowledge that, independently of this research, there are certain risks that pertain to all LLMs, as such models may generate outputs that, although plausible, are factually incorrect or nonsensical. Such hallucinations can misguide decision-making and propagate biases, especially in critical scenarios where accuracy is vital. Without proper safeguards, widespread LLM adoption could worsen these concerns. Thus, it is essential to develop mechanisms to mitigate hallucination risks, ensuring responsible and beneficial deployment of these powerful models before adopting the proposed model.
Acknowledgments
We are grateful to the Campus Super Computing Center at MBZUAI for supporting this work.
Appendix A Prompting Templates and Answer Extraction
Considering diverse LLMs and benchmarks adds challenges to prompting as no single uniform prompting approach best fits all LLMs (Sclar et al., 2023). Based on the insights from recent work on the appropriate usage of prompts (Sahoo et al., 2024) and from our own experiments, we make the following observations about the prompting trends: (1) For non-chat LLMs, few-shot chain-of-thought (COT) (Wei et al., 2022) prompting works better than zero-shot (Kojima et al., 2022) across both datasets, so we use 5 few-shot random examples obtained from the development set. The few-shot prompting leads to over 95% viable answers (except for llama2-7b-lm LLM which has the viability score of 83%) in generated solutions, where we consider an answer to be viable if it is represented by a single numeric/alphabetic string that can be extracted from the generated solution to compare with the reference answer. At the same time, a viable answer can be correct or incorrect. Viability is estimated using an automated script and is verified by manual inspection. (2) For chat LLMs, few-shot COT distracts the generator, which leads to unexpected outputs, so the zero-shot COT works best. To ensure correctness, we utilize different models’ chat templates from HuggingFace. [^1] https://huggingface.co/docs/transformers/en/chat_templating The viability of answer extraction for chat models is $\sim$92%. Examples of zero-shot and few-shot prompting are presented in Appendix Figure 4.
Split GSM8K MMLU
Train 6,816 13,757
Validation 359 285
Test 1,319 1,530
Table 4: Statistics on the datasets: For the MMLU dataset, we have swapped the officially released training and test splits to achieve a similar distribution as the GSM8K dataset.
LLMs Chat? Spec? #params
llama2-7b $\times$ $\times$ 7B
llama2-13b-chat $\checkmark$ $\times$ 13B
mistral-7b $\times$ $\times$ 7B
mistral-7b-it $\checkmark$ $\times$ 7B
gemma-7b $\times$ $\times$ 7B
gemma-7b-it $\checkmark$ $\times$ 7B
metamath-7b $\times$ $\checkmark$ 7B
Table 5: List of diverse LLMs selected for this study.
Spec: Specialized LLM
LLM Prompt Type GSM8K (prompt/sec) MMLU (prompt/sec)
llama2-7b 4.21 2.30
gemma-7b 7.10 3.00
Few-shot COT mistral-7b 3.70 1.10
metamath-7b 4.70 2.40
gemma-7b-it 0.70 1.00
Zero-shot COT llama2-13b-chat 1.80 4.80
mistral-7b-it 3.70 1.80
Table 6: Runtime statistics on various LLMs over 10 generations for each input query. The timings are recorded using a single A100 GPU. ‘sec’ denotes seconds, and COT denotes Chain-of-thought. For few-shot COT, we have considered 5 random examples from the development set.
Figure 4: Sample GSM8K zero-shot COT prompt for a chat (or instruction-tuned) LLMs, and MMLU few-shot COT prompt for non-chat LLMs.
The adapted prompting approaches used in our LLM queries are designed to instruct LLMs to specify that their final answers should be provided at the very end of each of their responses. We thus use a simple answer extraction policy of selecting the last mentioned numerical value (for GSM8K) and multiple-choice option (for MMLU) from the generated responses. Figure 4 in the Appendix shows a sample generation example. Responses failing to include any final answer are considered invalid (‘INVALID’) and counted as incorrect responses. For MMLU, we evaluate the extracted options directly against the annotated correct answers (‘A’, ‘B’, ‘C’, and ‘D’) from the dataset. For GSM8K, questions where the absolute difference between the ground truth and predicted numerical answers is less than $\epsilon=0.1$ are evaluated as solved correctly. This threshold was set to accommodate instances where model-generated real-valued answers differ slightly from the expected answers, e.g., due to rounding errors.
Figure 5: The distribution of different LLMs in the predicted subset of LLMs with SelectLLM algorithm for both GSM8K (top) and MMLU (bottom) datasets.
Figure 6: Distribution of the top-3 individual LLMs and proposed model responses to input queries for the test split of the GSM8K (left) and MMLU (right) datasets. The count in the rightmost column of each figure indicates the number of queries whose responses have been correctly answered by each LLM or the proposed model. The counts at the bottom denote the number of queries that have correct answers from one or more LLMs or the proposed model.
Appendix B Data Statistics and Selected LLMs
The statistics of the dataset and the list of considered diverse LLMs are presented in Table 5 and 5, respectively.
Appendix C LLM Latency Estimation
Runtime statistics of various LLMs are presented in Table 6.
Appendix D Sample Prompts
Prompting examples of few-shot COT and zero-shot COT are illustrated in Figure 4.
Appendix E Distribution of Different LLMs in Top-$s$ Subset Selected with SelectLLM
The details are shown in Figure 5.
Appendix F Query Response Distribution with Different Models
The details are presented in Figure 6.
Appendix G Implementation Details
Querying LLMs
We use the vLLM [^2] https://github.com/vllm-project/vllm package to query LLMs. All models were queried with a temperature of $0.8$ and a max token length of $2000$. Each question prompt was queried 10 times with different initialization seeds. We used a single NVIDIA A100 GPU for all runs. Querying each dataset once took approximately 1-2 hours.
MLC Training
We use the HuggingFace [^3] https://huggingface.co/ library for loading and tuning all pre-trained Transformer encoders in our experiments. Each model was trained for $10$ epochs, with an initial learning rate of $1e$-$6$, a warmup ratio of $0.1$, and class-balanced CrossEntropy loss. The training checkpoint with the lowest validation loss was selected for inference.
Appendix H Ablation Studies
The details are presented in Figure 7.
Figure 7: Ablation studies with top-$k$ and bottom-$k$ LLMs for GSM8K dataset (left). Similar observations are made on the MMLU dataset (right). For each LLM set, we have considered the values with an optimal $s$ for which the top-$s$ value is the highest.
Appendix I Importance of Domain: A Case Study with MMLU
The MMLU dataset comprises 57 subjects. In this analysis, we evaluate the performance of the proposed SelectLLM algorithm using the WeightedMaxConf policy, employing the best-performing individual LLM (gemma-7b-lm) on a subject-wise basis. Figure 8 illustrates that while the proposed model’s performance may be subpar for a few subjects, it demonstrates high performance for a significant portion of the subjects, indicating the effectiveness of the proposed model in general.
Figure 8: Subject-wise relative accuracy gain by SelectLLM with WeightedMaxConf policy over the performance of the best-performing individual LLM (gemma-7b-lm).